OpenStudy (anonymous):

MEDALS............................................................................................................ Find the volume of a triangular prism with the following dimensions Triangle base: 5 ft Triangle height: 3 ft Prism height: 7.5 ft

OpenStudy (anonymous):

ok the volume of a triangular prism is the area of the base(height)

OpenStudy (anonymous):

56 ft ^ 3 19 ft ^ 3 113 ft ^ 3 8 ft ^ 3

OpenStudy (anonymous):

the base of a triangular prism is a triangle...it's area can be calculated by 1/2(base)height

OpenStudy (anonymous):

so the area of the base of a triangular prism is 1/2(5)(3)=15/2

OpenStudy (anonymous):

therefore the volume of the prism=area of triangular base(height of prism)...plugging in area=15/2 and height of prism=7/5 feet you get...

OpenStudy (anonymous):

15/2(7.5)=56.25 ft^3---A

OpenStudy (anonymous):

ty